The absorbance of a solution of 106mg of a compound Butyraldehyde in 100 ml of 95% ethanol in a cell of 1 cm pathlength was found to be 0.28 calculate the molar absorptivity of the compound ?


1
Expert's answer
2021-04-11T23:51:08-0400

Molar Mass of Butyraldehyde = 72.11

= 0.106/72.11 = 0.00147 moles

Concentration = 0.00147/(0.095+0.106)

= 0.0073mol/L

= 0.28(1×0.0073)

= 0.002044 L mol-1 Cm-1